Two mechanics worked on a car. The first mechanic charged
$65
per hour, and the second mechanic charged
$100
per hour. The mechanics worked for a combined total of
25
hours, and together they charged a total of
$1800
. How long did each mechanic work?

Answer by jorel555(1290)

Let m be the 1st mechanic, and n be the 2nd. Then:
m+n=25
65m+100n=1800
100m+100n=2500
35m=700
m=20
n=5.
The 1st mechanic worked 20 hours; the 2nd one worked 5 hours. ☺☺☺☺